40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 773     Entry time: Thu Dec 29 01:11:32 2011
Author: Frank 
Type: DailyProgress 
Category: Seismic 
Subject: Accelerometer noise floor update 

found the old data from the seismometer measurements we did some time ago. Measurements agree more or less. Seismic noise floor measured some time ago was higher above 30Hz, but that was measured at the time where we had the construction going on and not during vacation time.

accelerometer_v2.png

accelerometer_v2.fig

ELOG V3.1.3-